OpenStudy (anonymous):

Charles got a raise from $5.00 an hour to $7.00 an hour. This was a raise of what percent? A. 2% B. 20% C. 40% D. 50%

OpenStudy (anonymous):

B ?

OpenStudy (lncognlto):

Um, not quite: 7.00 - 5.00 = 2.00 2.00/5.00 x 100 = ? That should give you the answer :)

OpenStudy (anonymous):

so A ?

OpenStudy (anonymous):

The answer is c

OpenStudy (lncognlto):

2/5 = 0.4 0.4 x 100 = 40 So then the answer would be C.